# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 3.4.2 Multi-Agent System Optimization
The optimization of agentic workflows focuses on finding the most effective, often static, structure of communication and task delegation for a given problem. Early research established important foundations, with studies like AutoFlow li2024autoflow demonstrating the automated creation of linear workflows from natural language, and GPTSwarm zhuge2024gptswarm proposing a unifying graph-based framework. Concurrently, other foundational work explored how agents could evolve by using symbolic learning to distill their interaction experiences into an explicit, interpretable set of logical rules to guide future decisions zhou2024symbolic . This abstraction of systems into tunable components—whether nodes, edges, or symbolic rules—was crucial. However, these early systems often lacked a formal method for efficiently navigating the vast space of possible configurations and interactions.
@@
The major breakthrough came when ADAS hu2024automated and AFlow zhang2024aflow formally defined this challenge as a search and optimization problem. ADAS set a theoretical vision by framing system design as a search through a Turing-complete space of code-based configurations. Building on this, AFlow made it practical by introducing reusable operators that represent common agentic patterns and by employing Monte Carlo Tree Search (MCTS) to efficiently navigate the enormous design space. Together, these works established a core methodology for treating agent system design as a tractable optimization problem, proving that automatically discovered workflows could outperform human-designed ones.
@@
Following this formalization, research rapidly diversified toward creating customized agent systems for each specific query. Two primary strategies emerged: search-based and learning-based generation. Search-based methods, such as MaAS zhang2025maas , create a "supernet" of potential architectures and then sample a specialized system from it. In parallel, learning-based methods train models to generate effective topologies directly. ScoreFlow wang2025scoreflow , for instance, trains a generator using a novel preference optimization method, while FlowReasoner gao2025flowreasoner uses reinforcement learning to train a meta-agent that constructs a bespoke workflow on the fly. This line of query-specific generation continues to be an active area of research ye2025masgpt ; ke2025maszero . Furthermore, it is important to note that this process is not limited to the topology alone; many of these frameworks also perform node-level optimization in tandem, such as co-optimizing prompts or selecting heterogeneous models as an integral part of the architectural generation process zhang2024aflow ; zhou2025maasprompt ; zhang2025evoflow .
@@
A key challenge for all search and learning methods is the computational cost of evaluating each potential workflow shang2025agentsquare . To address this, researchers have developed lightweight prediction models. Agentic Predictor trirat2025agentic is a prime example, training a model to accurately estimate a workflow’s performance based on its structural and semantic features without a full execution. By providing a fast and inexpensive evaluation proxy, these predictors significantly accelerate the optimization process, making the exploration of vast design spaces feasible zhang2025gnns .
@@
Distinct from optimizing a system’s explicit workflow structure, this line of research focuses on how multiple autonomous agents can co-evolve their internal behavioral policies through interaction. This approach enables emergent capabilities like coordination, task delegation, and beneficial competition. For instance, ReMA wan2025remalearningmetathinkllms uses multi-agent reinforcement learning (MARL) to collaboratively train a high-level meta-thinker and a low-level executor, significantly improving performance on reasoning benchmarks. Building on this, GiGPO feng2025groupingrouppolicyoptimizationllm enhances MARL training by aggregating trajectories to provide more precise credit assignment, boosting success rates on long-horizon tasks. To support this direction, platforms like MARTI liao2025marftmultiagentreinforcementfinetuning provide open-source infrastructure for orchestrating and scaling the training of these language-model collectives. Collectively, these studies underscore multi-agent reinforcement learning as a promising route for cultivating group-level competencies unattainable by individual agents alone.
@@ 4 When to Evolve
Intra-test-time self-evolution refers to adaptive processes that occur during task execution, where agents recognize their limitations on a specific problem and initiate targeted learning mechanisms to enhance their capabilities in real-time xi2024enhancing ; bi2024forest . This mode of evolution is characterized by its immediate coupling with the task at hand: the agent improves its problem-solving abilities for a specific problem encountered, creating a dynamic interplay between performance and adaptation.
Inter-test-time self-evolution refers to learning processes that occur between task completions, leveraging accumulated experiences to improve future performance. This category encompasses diverse methodological approaches: offline learning paradigms that extract knowledge from pre-collected datasets through iterative refinement zelikman2022star ; zelikman2024quiet , and online learning paradigms that continuously adapt based on streaming interaction data qi2024webrl ; qiu2025alita ; qian2024new ; wang2025mobile .
The implementation of self-evolution across these temporal phases leverages three fundamental learning paradigms in LLMs: in-context learning (ICL) dong2022survey ; min2021metaicl ; wies2023learnability , which adapts behavior through contextual examples without modifying parameters; supervised fine-tuning (SFT), which updates model weights through gradient-based optimization on labeled data devlin2018bert ; shen2024rethinking ; dong2023abilities ; and reinforcement learning (RL), which shapes behavior through reward-driven policy optimization kaelbling1996reinforcement ; sun2024llm ; zhang2025nemotron . While these learning paradigms remain conceptually consistent across temporal contexts, their instantiation differs in terms of data availability and learning objectives:
@@ 4.1 Intra-Test-Time Self-Evolution
Intra-test-time ICL methods leverage the model’s context window as a dynamic memory system for immediate adaptation without parameter modification. These approaches typically employ self-reflective mechanisms where agents analyze their own performance, generate verbal critiques or insights, and maintain these reflections in episodic memory buffers to guide subsequent decisions within the same task context shinn2023reflexion ; madaan2023self_refine . Some methods extend beyond simple reflection to include dynamic planning revision, where agents can modify their entire approach based on environmental feedback, switching between action execution and plan modification as needed. For instance, AdaPlanner sun2023adaplanner decomposes tasks into manageable sub-goals and predicts environmental feedback for each. During execution, its refiner component distinguishes between in-plan feedback (observations aligning with predictions) and out-of-plan feedback (deviating observations). For in-plan feedback, the refiner dynamically queries the LLM through a specialized ask_LLM() action to parse observations and extract pertinent information. For out-of-plan feedback, the refiner proactively revises the entire plan and resumes solving from an intermediate point, rather than restarting from scratch. This adaptive closed-loop framework eliminates the need for prior knowledge about feedback structures and enables more efficient decision-making. Similarly, TrustAgent hua2024trustagent employs rule-based plan revision during execution, modifying its approach based on language feedback to evolve toward safer planning strategies. These ICL methods demonstrate how test-time adaptation can achieve sophisticated behavioral modification without permanent model changes, maintaining flexibility while preserving the model’s general capabilities.
@@
Intra-test-time SFT represents a paradigm shift where models perform immediate self-modification through learned meta-adaptation strategies. Self-adaptive language modeling zweiger2025self exemplifies this approach by generating “self-edits”, which are meta-level instructions that can restructure information representations, specify optimization hyperparameters, or invoke tools for data augmentation and gradient computation. These self-edits trigger immediate supervised fine-tuning, resulting in persistent weight updates that adapt the model to the current task. The key innovation lies in the meta-learning phase, where reinforcement learning trains models to produce effective self-edits by using the downstream performance of the updated model as the reward signal, essentially teaching models how to teach themselves.
@@
Intra-test-time RL enables models to develop new capabilities on-demand when encountering problems beyond their current competence. LADDER simonds2025ladder demonstrates this through its test-time reinforcement learning (TTRL) mechanism: upon identifying a particularly challenging problem, the system generates a focused set of related problem variants and conducts intensive, targeted reinforcement learning specifically for that problem class. This approach transforms insurmountable challenges into learning opportunities, allowing models to expand their problem-solving repertoire during deployment rather than failing or providing suboptimal solutions. The method represents a form of just-in-time skill acquisition, where computational resources are invested precisely when and where they are needed most.
@@ 4.2 Inter-Test-Time Self-Evolution
Inter-test-time self-evolution represents the predominant learning process in autonomous agents, wherein adaptation occurs following task execution rather than during it. In this temporal mode, agents complete a given task, extract feedback signals, including explicit rewards gao2024designing , gradients amari1993backpropagation ; bottou2010large , and performance metrics ge2023openagi , and subsequently leverage this information to enhance their capabilities for future problem-solving. This retrospective learning process decouples task performance from capability improvement, allowing agents to consolidate experiences, identify patterns of success and failure, and systematically refine their behavioral policies without the computational constraints imposed by real-time task demands.
@@
Inter-test-time in-context learning has emerged as a widely adopted approach for agent self-improvement. This paradigm leverages execution results and feedback from previous tasks as contextual information for future problem-solving. Wang et al. wang2024agent demonstrate this principle by inducing workflows from agent action histories and incorporating them into the context for subsequent tasks. The field of in-context reinforcement learning (ICRL) moeini2025survey ; laskin2022context ; lee2023supervised extends this concept by maintaining histories of observations and actions within the agent’s context window. These methods exploit the hypothesis that pre-trained neural networks can implement implicit reinforcement learning algorithms within their forward pass, processing contextual information to adapt behavior without parameter updates kirsch2023towards . A defining characteristic of ICRL is in-context improvement: the phenomenon whereby agent performance progressively enhances as task-relevant information accumulates in the context, enabling sophisticated adaptation through attention mechanisms rather than gradient-based learning.
@@
Inter-test-time SFT chen2025sft methods establish a paradigm of iterative self-improvement through synthetic data generation and self-evaluation. SELF lu2023self pioneered meta-cognitive training, where models first acquire self-feedback and self-refinement capabilities, then iteratively generate responses to unlabeled instructions and enhance them through self-critique. STaR zelikman2022star and Quiet-STaR zelikman2024quiet focus on reasoning improvement through rationalization—models attempt problems, then generate explanations for correct answers they initially failed to solve, creating augmented training data that combines successful attempts with post-hoc reasoning. SiriuS zhao2025sirius extends this to sequential problem-solving, maintaining repositories of correct solutions while augmenting failures through multi-stage refinement involving feedback incorporation, regeneration, and rephrasing. These methods share a core insight: models can bootstrap their own improvement by learning to evaluate and enhance their outputs, creating high-quality training signals from initially imperfect attempts without extensive human supervision.
@@
Inter-test-time RL leverages unconstrained computational resources to optimize agents through extensive environmental interaction and sophisticated curriculum design. RAGEN wang2025ragen and DYSTIL wang2025dystildynamicstrategyinduction employ online reinforcement learning for multi-turn interactive tasks, continuously refining policies through on-policy learning in simulated dialogues. Learning Like Humans zhang2025learning introduces cognitive-inspired training with adaptive difficulty progression, combining on-policy exploration with off-policy efficiency and expert demonstrations to accelerate learning. Domain-specific applications demonstrate the versatility of inter-test-time RL: WebRL qi2024webrl develops web navigation agents through self-evolving curricula that automatically adjust task complexity based on performance, while DigiRL bai2024digirl enables device-control agents to master in-the-wild interactions through autonomous reinforcement learning. These approaches exploit the pre-deployment phase to engage in extensive trial-and-error learning, developing robust policies through thousands of interactions that would be impractical during real-time deployment.
@@ 5 How to Evolve
Reflexion shinn2023reflexion language internal ICL context test-time
AdaPlanner sun2023adaplanner language external + internal ICL context test-time
AgentS2 agashe2025agents2 language external ICL context test-time
@@
SELF lu2023self language external + internal SFT full params pre-test time + test-time
@@
SELF-REFINE madaan2023self_refine language internal ICL context test-time
@@
SCoRe kumar2024training numerical external RL full params pre-test time
PAG jiang2025pag numerical external RL full params pre-test time
@@
TextGrad yellamraju2024textgrad language external ICL context pre-test time / test-time
@@
SRSI simonds2025selfrewardingselfimproving language internal RL full params pre-test time
@@
Self-Train LM shafayat2025largereasoningmodelsselftrain numerical internal RL full params pre-test time
MM-UPT wei2025unsupervisedposttrainingmultimodalllm numerical internal RL full params pre-test time
@@
CoVo zhang2025consistentpathsleadtruth numerical internal RL full params pre-test time
SWE-agent du2025swedevevaluatingtrainingautonomous language external ICL context test-time
@@
SICA robeyns2025self numerical external ICL codebase(tools, workflows, prompts) test-time
@@
Feedback Friction jiang2025feedbackfriction language external ICL context test-time
USEagent applis2025useagent language external ICL context test-time
@@
DYSTIL wang2025dystildynamicstrategyinduction language + numerical external + internal SFT+RL full params pre-test time + test-time
OTC-PO wang2025otcpo numerical external RL full params pre-test time
@@
AUTORULE wang2025autorule language + numerical external + internal RL full params pre-test time
@@
EGSR zhang2025learning numerical external RL full params pre-test time
@@
LADDER simonds2025ladder numerical external RL full params pre-test time
@@
RAGEN wang2025ragen numerical external RL full params test-time
SPIRAL liu2025spiral numerical internal RL full params pre-test time
@@
ICRL Prompting song2025reward numerical external + internal RL full params test-time
@@
MATH-SHEPHERD wang2023math numerical external RL full params pre-test time
@@
AgentPRM choudhury2025process numerical external SFT+RL full params pre-test time
@@
Agent Q putta2024agentq numerical external RL full params pre-test time
@@
GiGPO feng2025group numerical external RL full params pre-test time
SPA-RL wang2025spa numerical external RL full params pre-test time
@@
Self-Instruct wang2022self language internal SFT full params pre-test time
WizardLM xu2024wizardlm language internal SFT full params pre-test time
OS-Genesis sun2024genesis numerical external SFT full params pre-test time
@@
UI-Genie xiao2025ui numerical external SFT partial params pre-test time
GUI-R1 luo2025gui numerical external SFT+RL full params pre-test time
InfiGUI-R1 liu2025infigui numerical external SFT+RL full params pre-test time
Voyager wang2023voyager language external ICL context test-time
SwiftSage lin2023swiftsage language external ICL context test-time
AutoWebGLM lai2024autowebglm language external SFT+RL full params pre-test time
DigiRL bai2024digirl language external RL partial params pre-test time
WebRL qi2024webrl language external SFT+RL full params pre-test time
@@
Let’s Verify Step-by-Step lightman2023let language external SFT full params pre-test time
@@
AlphaMath chen2024alphamath numerical external SFT full params pre-test time
@@
rStar-Math guan2025rstar numerical external SFT full params pre-test time
@@
DistRL wang2024distrl language external RL full params pre-test time + test-time
MobileGUI-RL shi2025mobileguirladvancingmobilegui language external RL full params pre-test time
@@
STaR zelikman2022star language + numerical internal SFT full params pre-test time
V-STaR hosseini2024v numerical external + internal SFT + RL partial params pre-test time
AdaSTaR koh2025adastar numerical internal SFT full params pre-test time
STIC deng2024enhancing language internal RL + SFT partial params pre-test time
GENIXER zhao2024genixer language external SFT full params pre-training
SiriuS zhao2025sirius language + numerical internal SFT full params pre-test time
SOFT tang2025bridging language internal SFT not specified pre-test time
RISE qu2024recursive language + numerical internal + external SFT full params pre-test time
IoE li2024confidence numerical internal / / test-time
@@
DGM zhang2025darwin numerical external ICL codebase (tools, workflows, prompts) test-time
EvoMAC hu2024self language external ICL team composition, workflow, prompts test-time
SPIN chen2024self language internal RL full params pre-test time
GENOME zhang2025nature numerical external Evolution Alg. partial params pre-test time
SPC chen2025spc numerical internal SFT+RL critic params pre-test time + test-time
Puppeteer dang2025multi numerical external RL planner policy pre-test time / between tasks
MedAgentSim almansoori2025self language external ICL context (knowledge base) test-time
STL mendes2025language language + numerical internal SFT value model pre-test time
MDTeamGPT chen2025self language external ICL context (knowledge base) test-time

### v4

@@ 3.4.2 Multi-Agent System Optimization
The optimization of agentic workflows focuses on finding the most effective, often static, structure of communication and task delegation for a given problem. Early research established important foundations, with studies like AutoFlow (Li et al., 2024d) demonstrating the automated creation of linear workflows from natural language, and GPTSwarm (Zhuge et al., 2024) proposing a unifying graph-based framework. Concurrently, other foundational work explored how agents could evolve by using symbolic learning to distill their interaction experiences into an explicit, interpretable set of logical rules to guide future decisions (Zhou et al., 2024b). This abstraction of systems into tunable components—whether nodes, edges, or symbolic rules—was crucial. However, these early systems often lacked a formal method for efficiently navigating the vast space of possible configurations and interactions.
@@
The major breakthrough came when ADAS (Hu et al., 2024c) and AFlow (Zhang et al., 2024c) formally defined this challenge as a search and optimization problem. ADAS set a theoretical vision by framing system design as a search through a Turing-complete space of code-based configurations. Building on this, AFlow made it practical by introducing reusable operators that represent common agentic patterns and by employing Monte Carlo Tree Search (MCTS) to efficiently navigate the enormous design space. Together, these works established a core methodology for treating agent system design as a tractable optimization problem, proving that automatically discovered workflows could outperform human-designed ones.
@@
Following this formalization, research rapidly diversified toward creating customized agent systems for each specific query. Two primary strategies emerged: search-based and learning-based generation. Search-based methods, such as MaAS (Zhang et al., 2025d), create a "supernet" of potential architectures and then sample a specialized system from it. In parallel, learning-based methods train models to generate effective topologies directly. ScoreFlow (Wang et al., 2025m), for instance, trains a generator using a novel preference optimization method, while FlowReasoner (Gao et al., 2025) uses reinforcement learning to train a meta-agent that constructs a bespoke workflow on the fly. This line of query-specific generation continues to be an active area of research (Ye and others, 2025; Ke et al., 2025). Furthermore, it is important to note that this process is not limited to the topology alone; many of these frameworks also perform node-level optimization in tandem, such as co-optimizing prompts or selecting heterogeneous models as an integral part of the architectural generation process (Zhang et al., 2024c; Zhou et al., 2025a; Zhang et al., 2025c).
@@
A key challenge for all search and learning methods is the computational cost of evaluating each potential workflow (Shang et al., 2025). To address this, researchers have developed lightweight prediction models. Agentic Predictor (Trirat et al., 2025) is a prime example, training a model to accurately estimate a workflow’s performance based on its structural and semantic features without a full execution. By providing a fast and inexpensive evaluation proxy, these predictors significantly accelerate the optimization process, making the exploration of vast design spaces feasible (Zhang et al., 2025s).
@@
Distinct from optimizing a system’s explicit workflow structure, this line of research focuses on how multiple autonomous agents can co-evolve their internal behavioral policies through interaction. This approach enables emergent capabilities like coordination, task delegation, and beneficial competition. For instance, ReMA (Wan et al., 2025) uses multi-agent reinforcement learning (MARL) to collaboratively train a high-level meta-thinker and a low-level executor, significantly improving performance on reasoning benchmarks. Building on this, GiGPO (Feng et al., 2025b) enhances MARL training by aggregating trajectories to provide more precise credit assignment, boosting success rates on long-horizon tasks. To support this direction, platforms like MARTI (Liao et al., 2025) provide open-source infrastructure for orchestrating and scaling the training of these language-model collectives. Collectively, these studies underscore multi-agent reinforcement learning as a promising route for cultivating group-level competencies unattainable by individual agents alone.
@@ 4 When to Evolve
Intra-test-time self-evolution refers to adaptive processes that occur during task execution, where agents recognize their limitations on a specific problem and initiate targeted learning mechanisms to enhance their capabilities in real-time (Xi et al., 2024; Bi et al., 2024). This mode of evolution is characterized by its immediate coupling with the task at hand: the agent improves its problem-solving abilities for a specific problem encountered, creating a dynamic interplay between performance and adaptation.
Inter-test-time self-evolution refers to learning processes that occur between task completions, leveraging accumulated experiences to improve future performance. This category encompasses diverse methodological approaches: offline learning paradigms that extract knowledge from pre-collected datasets through iterative refinement (Zelikman et al., 2022; 2024), and online learning paradigms that continuously adapt based on streaming interaction data (Qi et al., 2024; Qiu et al., 2025b; Qian et al., 2024; Wang et al., 2025n).
The implementation of self-evolution across these temporal phases leverages three fundamental learning paradigms in LLMs: in-context learning (ICL) (Dong et al., 2022; Min et al., 2021; Wies et al., 2023), which adapts behavior through contextual examples without modifying parameters; supervised fine-tuning (SFT), which updates model weights through gradient-based optimization on labeled data (Devlin et al., 2018; Shen, 2024; Dong et al., 2023); and reinforcement learning (RL), which shapes behavior through reward-driven policy optimization (Kaelbling et al., 1996; Sun et al., 2024a; Zhang et al., 2025n). While these learning paradigms remain conceptually consistent across temporal contexts, their instantiation differs in terms of data availability and learning objectives:
@@ 4.1 Intra-Test-Time Self-Evolution
Intra-test-time ICL methods leverage the model’s context window as a dynamic memory system for immediate adaptation without parameter modification. These approaches typically employ self-reflective mechanisms where agents analyze their own performance, generate verbal critiques or insights, and maintain these reflections in episodic memory buffers to guide subsequent decisions within the same task context (Shinn et al., 2023; Madaan et al., 2023a). Some methods extend beyond simple reflection to include dynamic planning revision, where agents can modify their entire approach based on environmental feedback, switching between action execution and plan modification as needed. For instance, AdaPlanner (Sun et al., 2023) decomposes tasks into manageable sub-goals and predicts environmental feedback for each. During execution, its refiner component distinguishes between in-plan feedback (observations aligning with predictions) and out-of-plan feedback (deviating observations). For in-plan feedback, the refiner dynamically queries the LLM through a specialized ask_LLM() action to parse observations and extract pertinent information. For out-of-plan feedback, the refiner proactively revises the entire plan and resumes solving from an intermediate point, rather than restarting from scratch. This adaptive closed-loop framework eliminates the need for prior knowledge about feedback structures and enables more efficient decision-making. Similarly, TrustAgent (Hua et al., 2024) employs rule-based plan revision during execution, modifying its approach based on language feedback to evolve toward safer planning strategies. These ICL methods demonstrate how test-time adaptation can achieve sophisticated behavioral modification without permanent model changes, maintaining flexibility while preserving the model’s general capabilities.
@@
Intra-test-time SFT represents a paradigm shift where models perform immediate self-modification through learned meta-adaptation strategies. Self-adaptive language modeling (Zweiger et al., 2025) exemplifies this approach by generating “self-edits”, which are meta-level instructions that can restructure information representations, specify optimization hyperparameters, or invoke tools for data augmentation and gradient computation. These self-edits trigger immediate supervised fine-tuning, resulting in persistent weight updates that adapt the model to the current task. The key innovation lies in the meta-learning phase, where reinforcement learning trains models to produce effective self-edits by using the downstream performance of the updated model as the reward signal, essentially teaching models how to teach themselves. Acikgoz et al. (2025) introduce a Test-Time Self-Improvement (TT-SI) framework that enables agents to adapt on-the-fly by first identifying uncertain test samples through self-awareness. For these challenging inputs, the agent then generates a single synthetic training example and performs a temporary, lightweight parameter update to improve its immediate performance before resetting its weights.
@@
Intra-test-time RL enables models to develop new capabilities on-demand when encountering problems beyond their current competence. LADDER (Simonds and Yoshiyama, 2025) demonstrates this through its test-time reinforcement learning (TTRL) mechanism: upon identifying a particularly challenging problem, the system generates a focused set of related problem variants and conducts intensive, targeted reinforcement learning specifically for that problem class. This approach transforms insurmountable challenges into learning opportunities, allowing models to expand their problem-solving repertoire during deployment rather than failing or providing suboptimal solutions. The method represents a form of just-in-time skill acquisition, where computational resources are invested precisely when and where they are needed most.
@@ 4.2 Inter-Test-Time Self-Evolution
Inter-test-time self-evolution represents the predominant learning process in autonomous agents, wherein adaptation occurs following task execution rather than during it. In this temporal mode, agents complete a given task, extract feedback signals, including explicit rewards (Gao et al., 2024), gradients (Amari, 1993; Bottou, 2010), and performance metrics (Ge et al., 2023), and subsequently leverage this information to enhance their capabilities for future problem-solving. This retrospective learning process decouples task performance from capability improvement, allowing agents to consolidate experiences, identify patterns of success and failure, and systematically refine their behavioral policies without the computational constraints imposed by real-time task demands.
@@
Inter-test-time in-context learning has emerged as a widely adopted approach for agent self-improvement. This paradigm leverages execution results and feedback from previous tasks as contextual information for future problem-solving. Wang et al. (Wang et al., 2024j) demonstrate this principle by inducing workflows from agent action histories and incorporating them into the context for subsequent tasks. The field of in-context reinforcement learning (ICRL) (Moeini et al., 2025; Laskin et al., 2022; Lee et al., 2023) extends this concept by maintaining histories of observations and actions within the agent’s context window. These methods exploit the hypothesis that pre-trained neural networks can implement implicit reinforcement learning algorithms within their forward pass, processing contextual information to adapt behavior without parameter updates (Kirsch et al., 2023). A defining characteristic of ICRL is in-context improvement: the phenomenon whereby agent performance progressively enhances as task-relevant information accumulates in the context, enabling sophisticated adaptation through attention mechanisms rather than gradient-based learning.
@@
Inter-test-time SFT (Chen et al., 2025b) methods establish a paradigm of iterative self-improvement through synthetic data generation and self-evaluation. SELF (Lu et al., 2023) pioneered meta-cognitive training, where models first acquire self-feedback and self-refinement capabilities, then iteratively generate responses to unlabeled instructions and enhance them through self-critique. STaR (Zelikman et al., 2022) and Quiet-STaR (Zelikman et al., 2024) focus on reasoning improvement through rationalization—models attempt problems, then generate explanations for correct answers they initially failed to solve, creating augmented training data that combines successful attempts with post-hoc reasoning. SiriuS (Zhao et al., 2025b) extends this to sequential problem-solving, maintaining repositories of correct solutions while augmenting failures through multi-stage refinement involving feedback incorporation, regeneration, and rephrasing. These methods share a core insight: models can bootstrap their own improvement by learning to evaluate and enhance their outputs, creating high-quality training signals from initially imperfect attempts without extensive human supervision. Recent frameworks such as ARIA (He et al., 2025a) further extend this paradigm by incorporating human-in-the-loop guidance into test-time adaptation, allowing agents to proactively identify knowledge gaps and request expert feedback.
@@
Inter-test-time RL leverages unconstrained computational resources to optimize agents through extensive environmental interaction and sophisticated curriculum design. RAGEN (Wang et al., 2025p) and DYSTIL (Wang et al., 2025c) employ online reinforcement learning for multi-turn interactive tasks, continuously refining policies through on-policy learning in simulated dialogues. Learning Like Humans (Zhang et al., 2025a) introduces cognitive-inspired training with adaptive difficulty progression, combining on-policy exploration with off-policy efficiency and expert demonstrations to accelerate learning. Domain-specific applications demonstrate the versatility of inter-test-time RL: WebRL (Qi et al., 2024) develops web navigation agents through self-evolving curricula that automatically adjust task complexity based on performance, while DigiRL (Bai et al., 2024) enables device-control agents to master in-the-wild interactions through autonomous reinforcement learning. These approaches exploit the pre-deployment phase to engage in extensive trial-and-error learning, developing robust policies through thousands of interactions that would be impractical during real-time deployment.
@@ 5 How to Evolve
Reflexion(Shinn et al., 2023) language internal ICL context test-time
AdaPlanner(Sun et al., 2023) language external + internal ICL context test-time
AgentS2(Agashe et al., 2025) language external ICL context test-time
@@
SELF(Lu et al., 2023) language external + internal SFT full params pre-test time + test-time
@@
SELF-REFINE(Madaan et al., 2023a) language internal ICL context test-time
@@
SCoRe(Kumar et al., 2024) numerical external RL full params pre-test time
PAG(Jiang et al., 2025b) numerical external RL full params pre-test time
@@
TextGrad(Yellamraju and others, 2024) language external ICL context pre-test time / test-time
@@
SRSI(Simonds et al., 2025) language internal RL full params pre-test time
@@
Self-Train LM(Shafayat et al., 2025) numerical internal RL full params pre-test time
MM-UPT(Wei et al., 2025c) numerical internal RL full params pre-test time
@@
CoVo(Zhang et al., 2025j) numerical internal RL full params pre-test time
SWE-agent(Du et al., 2025) language external ICL context test-time
@@
SICA(Robeyns et al., 2025a) numerical external ICL codebase(tools, workflows, prompts) test-time
@@
Feedback Friction(Jiang et al., 2025a) language external ICL context test-time
USEagent(Applis et al., 2025) language external ICL context test-time
@@
DYSTIL(Wang et al., 2025c) language + numerical external + internal SFT+RL full params pre-test time + test-time
OTC-PO(Wang et al., 2025g) numerical external RL full params pre-test time
@@
AUTORULE(Wang and Xiong, 2025) language + numerical external + internal RL full params pre-test time
@@
EGSR(Zhang et al., 2025a) numerical external RL full params pre-test time
@@
LADDER(Simonds and Yoshiyama, 2025) numerical external RL full params pre-test time
@@
RAGEN(Wang et al., 2025p) numerical external RL full params test-time
SPIRAL(Liu et al., 2025d) numerical internal RL full params pre-test time
@@
ICRL Prompting(Song et al., 2025) numerical external + internal RL full params test-time
@@
MATH-SHEPHERD(Wang et al., 2023b) numerical external RL full params pre-test time
@@
AgentPRM(Choudhury, 2025) numerical external SFT+RL full params pre-test time
@@
Agent Q(Putta et al., 2024) numerical external RL full params pre-test time
@@
GiGPO(Feng et al., 2025a) numerical external RL full params pre-test time
SPA-RL(Wang et al., 2025d) numerical external RL full params pre-test time
@@
Self-Instruct(Wang et al., 2022) language internal SFT full params pre-test time
WizardLM(Xu et al., 2024a) language internal SFT full params pre-test time
OS-Genesis(Sun et al., 2024b) numerical external SFT full params pre-test time
@@
UI-Genie(Xiao et al., 2025a) numerical external SFT partial params pre-test time
GUI-R1(Luo et al., 2025b) numerical external SFT+RL full params pre-test time
InfiGUI-R1(Liu et al., 2025e) numerical external SFT+RL full params pre-test time
Voyager(Wang et al., 2023a) language external ICL context test-time
SwiftSage(Lin et al., 2023) language external ICL context test-time
AutoWebGLM(Lai et al., 2024) language external SFT+RL full params pre-test time
DigiRL(Bai et al., 2024) language external RL partial params pre-test time
WebRL(Qi et al., 2024) language external SFT+RL full params pre-test time
@@
Let’s Verify Step-by-Step(Lightman et al., 2023) language external SFT full params pre-test time
@@
AlphaMath(Chen et al., 2024a) numerical external SFT full params pre-test time
@@
rStar-Math(Guan et al., 2025) numerical external SFT full params pre-test time
@@
DistRL(Wang et al., 2024g) language external RL full params pre-test time + test-time
MobileGUI-RL(Shi et al., 2025b) language external RL full params pre-test time
@@
STaR(Zelikman et al., 2022) language + numerical internal SFT full params pre-test time
V-STaR(Hosseini et al., 2024) numerical external + internal SFT + RL partial params pre-test time
AdaSTaR(Koh et al., 2025) numerical internal SFT full params pre-test time
STIC(Deng et al., 2024) language internal RL + SFT partial params pre-test time
GENIXER(Zhao et al., 2024c) language external SFT full params pre-training
SiriuS(Zhao et al., 2025b) language + numerical internal SFT full params pre-test time
SOFT(Tang et al., 2025) language internal SFT not specified pre-test time
RISE(Qu et al., 2024) language + numerical internal + external SFT full params pre-test time
IoE(Li et al., 2024b) numerical internal / / test-time
@@
DGM(Zhang et al., 2025h) numerical external ICL codebase (tools, workflows, prompts) test-time
EvoMAC(Hu et al., 2024d) language external ICL team composition, workflow, prompts test-time
SPIN(Chen et al., 2024f) language internal RL full params pre-test time
GENOME(Zhang et al., 2025r) numerical external Evolution Alg. partial params pre-test time
SPC(Chen et al., 2025c) numerical internal SFT+RL critic params pre-test time + test-time
Puppeteer(Dang et al., 2025) numerical external RL planner policy pre-test time / between tasks
MedAgentSim(Almansoori et al., 2025a) language external ICL context (knowledge base) test-time
STL(Mendes and Ritter, 2025) language + numerical internal SFT value model pre-test time
MDTeamGPT(Chen et al., 2025e) language external ICL context (knowledge base) test-time